# Suppression of the Mott insulating phase in the particle-hole asymmetric Hubbard model

### v1

@@ 1 Introduction
The MIT has been extensively studied in the literature, mostly using the dynamical mean field theory (DMFT) approach [15, 16] in models with particle-hole symmetry (PHS). In this case, it is well established that, below a critical temperature, there is a first-order transition line in the temperature ($T$) versus interaction ($U$) phase diagram ending in the critical point $(U_{c},T_{c})$. For $T<T_{c}$, the system is metallic for $U<U_{c_{1}}(T)$ and insulating for $U>U_{c_{2}}(T)$ with a metastable coexistence region for values of $U$ between these two values [15, 17, 18, 19, 20, 21]. The two spinodal lines defined by $U_{c_{1}}(T)$ and $U_{c_{2}}(T)$ meet at the critical point $(U_{c},T_{c})$. For $T>T_{c}$, the transition becomes a crossover [11, 14, 22].
@@
Although less studied, the particle-hole asymmetric (PHA) Hubbard model is of particular interest since that is the prevalent situation for the band structure [17, 25, 26] of most materials. Particle-hole asymmetry can be implemented in the Hamiltonian model by including next-nearest neighbor hopping matrix elements at half-filling or by explicitly changing the onsite energy [4, 8, 14, 26, 27, 28, 29].
In this work we revisit this problem by studying the Mott metal-insulator transition (MIT) in the particle-hole asymmetric Hubbard model in the Bethe lattice. As argued above, although the MIT in the Hubbard model has been extensively studied, most of the studies focus on the particle-hole symmetric (PHS) case as this naturally renders the system to be in half-filling (that corresponds to an occupation number $n=1$) independently of the interaction strength $U$. A crucial fact is that the MIT can occur away from PHS and this will be marked by $n\rightarrow 1$ at the transition. In other words, at the insulating side of the MIT, the system will be both particle-hole asymmetric (PHA) and at half-filling at the same time.
While earlier studies have addressed the transition away from PHS at zero temperature [30], here we show that the PHA system has a rich phase diagram at non-zero temperatures. As such, our focus here is on the fate of the co-existence region away from PHS. In fact, our results show that this region is suppressed . This is characterized also by a lower value of the maximum temperature $T_{c}$ for which a first-order transition between the metallic and insulating phases takes place. For $T>T_{c}$, the transition is second-order, such that $T_{c}$ marks the “end” of the first-order transition line [23].
Although similar decreases in $T_{c}$ have been seen in doped [31] and disordered systems [32, 33], here we provide a systematic analysis of these effects in a “clean” system in which the on-site energy can be independently varied (say, by an external gate-voltage). As such, our approach can be extended to systems in which the density $n$ is not fixed a priori and can vary across the MI transition.
More importantly, in the realistic case of systems away from PHS, we show that the charge density can be a reliable marker of the true MIT transition for larger temperatures, introducing the interesting prospect of experimentally characterizing the MIT by the electron (or hole) densities. This allows one to characterize the transition by the charge density $n$ rather than by the zero-frequency spectral density $\rho(0)$, the usual parameter considered in theoretical works.
@@ 3.1 Signatures of the MIT from the DMFT results
Refer to captionRefer to caption
Figure 1: Occupation $n$ and zero-frequency spectral density $\rho(0)$ versus $U$ for the metal-insulator M $\rightarrow$ I and the insulator-metal I $\rightarrow$ M paths for temperatures below and above $T_{c}$.
@@ 3.2 Phase diagrams away from PHS
From these considerations, one can then represent the phase diagram of the system by color plots for $\rho_{\rm avg}(0)$ as a function of $T$ and $U$. These are shown in Figure 2 (a)-(d) for different values of the onsite energy $\epsilon_{d}$. These define $T$ vs $U$ phase diagrams showing the insulating (blue), metallic (red) and co-existence (green) regions. As expected [15, 24], the difference between $U_{c_{2}}$ and $U_{c_{1}}$ decreases as $T$ increases, such that the co-existence regions narrows to a single point at a critical temperature $T_{c}$ at which $U_{c_{2}}=U_{c_{1}}=U_{c}$
Refer to captionRefer to caption
@@
Interestingly, the value of $T_{c}$ (and the overall area of the co-existence region) decreases as the system is moved away from the particle-hole symmetric point ($\epsilon_{d}=0.5\,U$). This is illustrated in Fig. 2(e), which shows a sharp decrease in $T_{c}$ as $\epsilon_{d}$ is reduced from $0.5U$ (PHS point) to $0.44U$  [^2] Numerical instabilities in the NCA calculations prevented us from exploring the regime of temperatures lower than $T\sim 0.002$, which restricts our analysis to the range $0.44U\lesssim\epsilon_{d}\leq 0.5U$. In addition, the corresponding value of $U_{c}$ increases as $\epsilon_{d}$ is reduced in the same range, as depicted in Fig. 2(f).
Alternatively, phase diagrams can be visualized by plotting color maps of $\rho_{\rm avg}(0)$ in the $U\times\epsilon_{d}$ plane for a given temperature $T$. This is illustrated for $T=0.002$ in Figs 3(a) and (b) and for $T=0.004$ by Figs. 3(c) and (d). These plots show more clearly the narrowing of the coexistence region, corresponding to the green area in Figs. 3(a) and (c), and the red region in Figs. 3(b) and (d), as the system is moved away from the PHS point. These results indicate that the first-order transition line between the metallic and insulating phases is better defined when the system is away from PHS.
@@ 3.3 Signatures of the MIT in the charge density
An important point regarding the phase diagrams away from the PHS is that different phases in the system can be characterized not only the the zero-frequency spectral density $\rho(0)$ but also by the charge density $n$ since the insulating region is characterized by half-filling, namely, $n=1$. As such, the $T$ vs $U$ phase diagrams can be plotted by looking at $n$ alone. In an analogy with $\rho_{\rm avg}(0)$ and $\Delta\rho(0)$ defined in Eq. (3.1), it is useful to define the following quantities related to the charge density:
@@ 4 Summary and conclusions
Our results show that, in contrast with the well-studied particle-hole symmetric case, the PHA model exhibits a notably complex $U\times T$ phase diagram showing reduced value of the critical temperature $T_{c}$ as well as a narrower co-existence region between metallic and insulating phases at low temperature. This is illustrated in the $U\times T$ and $U\times\epsilon_{d}$ phase diagrams, indicating the “shrinking” co-existence region with increasing temperature and decreasing symmetry. The plots indicate that the transition between metallic and insulating phases is better defined when the system is away from PHS.
@@
Moreover, a key contribution of our work is the identification of the charge density $n$ as a robust marker for the MIT transition in systems where particle-hole symmetry is broken. Unlike traditional approaches that rely on the zero-frequency spectral density, the charge density provides a more accessible and potentially experimentally verifiable parameter, especially in high-temperature regimes.
In summary, our study highlights the significant impact of breaking particle-hole symmetry on the MIT in the Hubbard model. The results underscore the need to consider symmetry effects in theoretical predictions and experimental observations of phase transitions in strongly correlated electron systems.

### v2

@@ 1 Introduction
The MIT has been extensively studied in the literature, mostly using the dynamical mean field theory (DMFT) approach [15, 16] in models with particle-hole symmetry (PHS). In this case, it is well established that, below a critical temperature, there is a first-order transition line in the temperature ($T$) versus interaction ($U$) phase diagram ending in the critical point $(U_{c},T_{c})$. For $T<T_{c}$, the system is metallic for $U<U_{c_{1}}(T)$ and insulating for $U>U_{c_{2}}(T)$ with a coexistence region (where one of the phases is thermodynamically stable while the other is only metastable) for values of $U$ between these two values [15, 17, 18, 19, 20, 21]. The two spinodal lines defined by $U_{c_{1}}(T)$ and $U_{c_{2}}(T)$ meet at the critical point $(U_{c},T_{c})$. For $T>T_{c}$, the transition becomes a crossover [11, 14, 22].
@@
Although less studied, the particle-hole asymmetric (PHA) Hubbard model is of particular interest since that is the prevalent situation for the band structure [17, 25, 26] of most materials. Particle-hole asymmetry can be implemented in the Hamiltonian model by including next-nearest neighbor hopping matrix elements at half-filling or by explicitly changing the onsite energy [4, 8, 14, 26, 27, 28, 29]. As argued above, although the MIT in the Hubbard model has been extensively studied, most of the studies focus on the particle-hole symmetric (PHS) case as this naturally renders the system to be in half-filling (that corresponds to an occupation number $n=1$) independently of the interaction strength $U$. A crucial fact is that the MIT can occur away from PHS and this will be marked by $n\rightarrow 1$ at the transition [27, 29]. In other words, at the insulating side of the MIT, the system will be both particle-hole asymmetric (PHA) and at half-filling at the same time.
In this work we revisit this problem by studying the Mott metal-insulator transition (MIT) in the particle-hole asymmetric Hubbard model in the Bethe lattice. While earlier studies have addressed the transition away from PHS at zero [27, 30] or very low temperatures [29], here we show that the PHA system has a rich phase diagram at non-zero temperatures. As such, our focus here is on the fate of the co-existence region away from PHS at larger temperatures. In fact, our results show that this region is suppressed as the system moves away from PHS, a fact which has not been properly documented in previous works.
This behavior is characterized by a lower value of the maximum temperature $T_{c}$ for which a first-order transition between the metallic and insulating phases takes place. For $T>T_{c}$, the transition is second-order, such that $T_{c}$ marks the “end” of the first-order transition line [23]. As such, this critical temperature $T_{c}$ decreases as the system moves away from PHS. Although similar decreases in $T_{c}$ have been seen in doped [31] and disordered systems [32, 33], here we provide a systematic analysis of these effects in a “clean” system in which the on-site energy can be independently varied (say, by an external gate-voltage). As such, our approach can be extended to systems in which the density $n$ is not fixed a priori and can vary across the MI transition.
More importantly, in the realistic case of systems away from PHS, we show that the charge density can be a reliable marker of the true MIT transition not only at $T\sim 0$ [29] but also for larger temperatures, introducing the interesting prospect of experimentally characterizing the MIT by the electron (or hole) densities. This allows one to characterize the transition by the charge density $n$ rather than by the zero-frequency spectral density $\rho(0)$, the usual parameter considered in theoretical works.
@@ 3.1 Signatures of the MIT from the DMFT results
Figure 1: Occupation $n$ and zero-frequency spectral density $\rho(0)$ versus $U$ for the metal-insulator M $\rightarrow$ I and the insulator-metal I $\rightarrow$ M paths for temperatures below and above $T_{c}$. Both paths coincide for $T>T_{c}$.
@@ 3.2 Phase diagrams away from PHS
From these considerations, one can then represent the phase diagram of the system by color plots for $\rho_{\rm avg}(0)$ as a function of $T$ and $U$. These are shown in Figure 2 (a)-(d) for different values of the onsite energy $\epsilon_{d}$. These define $T$ versus $U$ phase diagrams showing the insulating (blue), metallic (red) and co-existence (green) regions. As expected [15, 24], the difference between $U_{c_{2}}$ and $U_{c_{1}}$ decreases as $T$ increases, such that the co-existence regions narrows to a single point at a critical temperature $T_{c}$ at which $U_{c_{2}}=U_{c_{1}}=U_{c}$
@@
Interestingly, the value of $T_{c}$ (and the overall area of the coexistence region) decreases as the system is moved away from the particle-hole symmetric point ($\epsilon_{d}=0.5\,U$). This is illustrated in Fig. 2(e), which shows a sharp decrease in $T_{c}$ as $\epsilon_{d}$ is reduced from $0.5U$ (PHS point) to $0.44U$  [^2] Numerical instabilities in the NCA calculations prevented us from exploring the regime of temperatures lower than $T\sim 0.002$, which restricts our analysis to the range $0.44U<\sim\epsilon_{d}\leq 0.5U$ In addition, the corresponding value of $U_{c}$ increases as $\epsilon_{d}$ is reduced in the same range, as depicted in Fig. 2(f). In fact, the $U_{c}$ vs $\epsilon_{d}$ curve is very well approximated by a quadratic scaling of the form $U_{c}(\epsilon_{d})-U_{c0}\sim(\epsilon_{d}-U/2)^{2}$ (solid line in Fig. 2(f)), where $U_{c0}$ is the minimal value at the particle-hole symmetry point $\epsilon_{d}=U/2$. This excellent agreement with a quadratic scaling is consistent with analytical expressions relating $U_{c}$ and the chemical potential in the context of the doping-driven Mott transition [29, 30].
Alternatively, phase diagrams can be visualized by plotting color maps of $\rho_{\rm avg}(0)$ and $\Delta\rho(0)$, defined in Eq. (5), in the $U\times\epsilon_{d}$ plane for a given temperature $T$. This is illustrated for $T=0.002$ in Figs. 3(a) and (b) and for $T=0.004$ by Figs. 3(c) and (d). We note that both $\rho_{\rm avg}(0)$ and $\Delta\rho(0)$ are symmetric with respect to the PHS value $\epsilon_{d}=0.5U$ for all temperatures. Consequently, the curves $U_{c}(\epsilon_{d})$ and $T_{c}(\epsilon_{d})$ shown in Fig. 2 are also symmetric with respect to $\epsilon_{d}=0.5U$. These plots show more clearly the narrowing of the co-existence region, corresponding to the green area in Figs. 3(a) and (c), and the red region in Figs. 3(b) and (d), as the system is moved away from the PHS point.
@@ 3.3 Signatures of the MIT in the charge density
An important point regarding the phase diagrams away from the PHS is that different phases in the system can be characterized not only by the zero-frequency spectral density $\rho(0)$ but also by the charge density $n$ since the insulating region is characterized by half-filling, namely, $n=1$. As such, the $T$ versus $U$ phase diagrams can be plotted by looking at $n$ alone. In an analogy with $\rho_{\rm avg}(0)$ and $\Delta\rho(0)$ defined in Eq. (5), it is useful to define the following quantities related to the charge density:
@@ 4 Summary and conclusions
Our results show that, in contrast with the well-studied particle-hole symmetric case, the PHA model exhibits a notably complex $U\times T$ phase diagram showing reduced value of the critical temperature $T_{c}$ as well as a narrower co-existence region between metallic and insulating phases at low temperature. This is illustrated in the $U\times T$ and $U\times\epsilon_{d}$ phase diagrams, indicating the “shrinking” co-existence region with increasing temperature and/or breaking the particle-hole symmetry. The plots indicate that the transition between metallic and insulating phases is better defined when the system is away from PHS.
@@
Moreover, a key contribution of our work is the identification of the charge density $n$ as a robust marker for the MIT transition in systems where particle-hole symmetry is broken since the transition will be marked by $n\rightarrow 1$. We notice that this behavior should occur in realistic material systems where PHS is broken by an asymmetric density of states, which can be schematically modeled, for instance, by a Bethe lattice incorporating second-neighbor hopping [56]. Unlike traditional approaches that rely on the zero-frequency spectral density, the charge density provides a more accessible and potentially experimentally verifiable parameter, especially in high-temperature regimes.
In summary, our study highlights the significant impact of breaking particle-hole symmetry on the MIT in the Hubbard model at finite temperatures. The results underscore the need to consider symmetry effects in theoretical predictions and experimental observations of phase transitions in strongly correlated electron systems.
5 Acknowledgements